OpenStudy (anonymous):

Determine whether the graph of y = x2 − 4x + 1 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (2, -3) Maximum; (2, -3) Minimum; (-3, 2) Maximum; (-3, 2)

OpenStudy (precal):

because x^2 is +1x^2 it is concave up, therefore it has a min value

OpenStudy (anonymous):

no wonder students get so damned confused the minimum value of a function is a NUMBER not an ordered pair they are asking for the vertex of the parabola

OpenStudy (anonymous):

first coordinate of the vertex is always \[-\frac{b}{2a}\] in your case it is \[-\frac{-4}{2\times 1}=2\]

OpenStudy (anonymous):

so x=2 but then what is the other number?

OpenStudy (anonymous):

whatever you get for \(y\) when you replace \(x\) by 2

OpenStudy (precal):

 perhaps the wording of the problem could be better yes that min is called the vertex just like satellite73 stated earlier

OpenStudy (anonymous):

don't forget if you know the first coordinate, you can find the second by substitution

OpenStudy (anonymous):

so then 2,-3

OpenStudy (precal):

yes and don't forget it is a min